Chapter 12. Chi-Square Test Saved Help Save & Exit Su Check my w Given the following contingency table, conduct a test for InCalculate the test statistic.
Answer #1

The table for observed frequency is ,

Variable A
Variable B 1 2 Total
1 36 31 (A1) = 67
2 35 51 (A2) = 86
Total (B1) = 71 (B2) = 82 N=153

The table for expected frequency is ,

(A)(B) ; i=1,2 and j=1,2

Variable A
Variable B 1 2
1 31.0915 35.9085
2 39.9085 46.0915

The table for ei is ,

Variable A
Variable B 1 2
1 41.68341 26.76247
2 30.69522 56.43123

Therefore , the test statistic is ,

-N = 155.5723-153 = 2.5723